OpenStudy (anonymous):

Which of the following has the same terminal side as -11pi/6? a. 30 degrees b. 150 degrees c. 210 degrees d. 330 degrees and why?

OpenStudy (anonymous):

pi/6 is 180/6 = 30 11 pi/6 is 11*30 = 330 Terminal side is 360 -330 = 30 degrees Answer is a

OpenStudy (anonymous):

So to find the terminal side, does that mean we always have to subtract it from 360?